OpenStudy (anonymous):

Solve the system x - y = 3 and 3x - 8y = 9. What is the value of y? A. 3 B. 0 C. 6 D. -2

OpenStudy (anonymous):

rearrange x-y=3 to be x=3+y the substitute into the second equation 3(3+y) -8y=9 9+3y-8y=9 -5y=0 y=0

OpenStudy (swag):

b